OpenStudy (anonymous):

How do I calculate the angles of a parallelogram if the acute angel 30'

OpenStudy (gudden):

do u mean that one of the given angle is 30 degrees ??

OpenStudy (anonymous):

yes

OpenStudy (gudden):

okay then as in a parallelogram... oposite angles are equal therefore .. two of the angles of the parallelogram would be 30 degree... now let the unknown angle be x... and since in a parallelogram opposite angles are equal and all the angles add upto 360 therefore 2x + 60 = 360 2x = 300 x = 150 degree.. so.. the angles of the parallelogram are... 30 degree , 150 degree , 30 degree and 150 degree...

OpenStudy (gudden):

so do u want me to explain it to u in more easy way ??

OpenStudy (anonymous):

yes plz

OpenStudy (gudden):

alright then..

OpenStudy (gudden):



OpenStudy (gudden):

i have drawn a parallelogram with one angle as x... and since in a parallelogram opposite sides are equal therefore u could see the labelling so done.... is it clear till here ??

OpenStudy (anonymous):

yup

OpenStudy (gudden):

so do u know what is the sum of the internal angles of a quadrilateral?

OpenStudy (anonymous):

amm nope

OpenStudy (gudden):

anywayz... now put the sum of the angles of this parallelogram as 360 i.e. x + x+30 + 30 = 360 => 2x + 60 = 360 taking 60 to the Right Hand side.. then it gives us 2x = 360 - 60 i.e. 2x = 300 now x = 300/2 = 150 and hence u get ur unknown angle..:D

OpenStudy (anonymous):

ok so the obtuse angles are 150

OpenStudy (gudden):

yes...:D...

OpenStudy (anonymous):

oo ok thanks a lot realy c:

OpenStudy (gudden):

mention not...:D. it feels great when we explain someone something...:D

OpenStudy (anonymous):

lol ya